# On the reproducibility of discrete-event simulation studies in health research: an empirical study using open models

### v1

@@ Abstract
Article word count: 6675
@@ 1.3 Healthcare simulation models
This study focuses on healthcare simulation models, specifically DES. DES has a wide range of applications, including healthcare operations and system design, medical decision-making, and infectious disease modelling.5 It is the most commonly used simulation method in healthcare.39, 40, 5, 41
In recent years, efforts have been made to enhance the sharing and reuse of healthcare simulation models. The Strengthening The Reporting of Empirical Simulation Studies-DES (STRESS-DES) guidelines from Monks et al. (2019)2 were developed to facilitate replication by providing technical details necessary to recreate a model. Zhang et al. (2020)42 developed a “generic reporting checklist” focused on model quality, derived from the International Society for Pharmacoeconomics and Outcomes Research (ISPOR)–Society for Medical Decision Making (SMDM) Modelling Good Research Practices Task Force reports. In 2024, the Sharing Tools and Artefacts for Reusable Simulations (STARS) framework was introduced to provide guidance on which artefacts should be shared in repositories to facilitate model reuse.43
These guidelines and frameworks (fully detailed in Appendix A) were developed because many healthcare DES models are not shared in ways that facilitate understanding, use, and reproducibility. A 2023 review44 of 564 healthcare DES papers (2019-2022) revealed only 8.3% shared model code, although this was improving over time. Among shared models, only 60% included a README, with just 32% explaining how to run the model. Less than half provided an open licence, and only 10.6% archived their models. Across all papers, only 12.8% used reporting guidelines.44
@@
This study aims to assess the reproducibility of published healthcare DES models by examining a sample of eight studies. It will identify factors that facilitate or hinder each reproduction and will evaluate adherence to relevant frameworks, badges and reporting guidelines, to understand how these influence reproducibility. By understanding which elements are most important when reporting results in articles and sharing code, the study seeks to identify key factors that can improve reproducibility in healthcare DES models. While not all studies may have been designed with computational reproducibility in mind, it is increasingly recognised as an important consideration for any research involving code,10 as it helps verify the integrity of findings. Adopting basic practices to improve reproducibility can enhance scientific rigour and is achievable without requiring advanced technical skills.
@@ 2 Methods
The study protocol (as summarised in Figure 2 and described below) was pre-registered on 20 June 2024.45 It was informed by prior reproduction studies,17, 13, 46, 47, 23 and refined through a pilot on a Python DES model by Allen et al. (2020).48
@@ 2.1 Sample
This study included eight models: seven from a 2018-2022 review44 and one from an additional search for older published models. Models were developed in Python 49 or R50, the most popular Free and Open Source Software (FOSS) languages for healthcare DES.44 Models were selected to ensure diversity across a range of factors including the health focus (e.g. healthcare condition, specific system), geographical context, and model complexity - as in Table 2.1.
The protocol initially planned to include six studies, but this was later expanded to eight during the research process. This allowed the inclusion of another “medium-sized” model 51 that was felt to better represent typical operational research applications of DES in health compared to some larger-scale models. An older model52 was also included, enabling consideration of reproducibility challenges from an older code base (approximately a decade ago), as opposed to more recent studies (with all others published within the last five years).
@@ 2.2 Reproduction
Scope. At least two team members reviewed the article and agreed on which experimental results (e.g. tables, figures, results described in the text) to reproduce. The repository was then archived on Zenodo.62
@@ 2.3 Evaluation
Articles were evaluated using two DES reporting guidelines: STRESS-DES from Monks et al. (2019),2 and the “generic reporting checklist” from Zhang et al. (2020).42 Repositories were evaluated against the STARS framework43 and criteria for journal badges from: ACM,26 National Information Standards Organisation (NISO),63 Centre for Open Science (COS),27 Institute of Electrical and Electronics Engineers (IEEE),28 and Psychological Science.64, 32 The evaluation was performed by one researcher, with consensus from a second for uncertain or unmet criteria. For reference, full guidelines and badge criteria are provided in Appendix A.
@@ 2.4 Synthesis and consolidation
Results from the reproduction, evaluation, and reflections on the facilitators and barriers encountered were shared via a Quarto website65 hosted on GitHub Pages.66 The repository was organised into a research compendium,67 with a README, conventional file organisation (e.g. data, methods, outputs), tests, and a Docker environment.68 A second researcher verified execution and reproducibility by running the model and/or tests (locally and via Docker). The final repository was archived on Zenodo,62 and results were shared with the original authors.
@@ 2.5 Citations and visualisations
Studies are not referenced by name in the results, to keep the emphasis on overarching trends rather than individual reproduction success. For transparency, full results are available, with links to the reproduction and original study repositories provided in Appendix C. For this article, plots were created using Python 3.10.14,49 Plotly 5.23.0,69 Matplotlib 3.9.270 and pandas 2.2.2,71, 72 with a full list of dependencies in the research repository.73 Other visualisations were created using Inkscape 1.3.2 (https://inkscape.org/) and Sketchpad v2022.2.21.0 (https://sketch.io/sketchpad/).
@@ 3.2 Evaluation of the repository
The repositories were also evaluated against journal badge criteria, with results for the ACM badges summarised here (Table 3.2) and the full evaluation in Appendix B. Only one repository met the standards for the “Artefacts Available” badge. None satisfied “Artefacts Functional” requirements, typically due to incomplete code (missing scenarios, outputs, etc.). Assuming that reproductions needed to be within a reasonable time and with minimal troubleshooting, only one study was eligible for “Results Reproduced”.
@@ 3.3 Evaluation of the article
Studies were evaluated against STRESS-DES2 and the generic reporting checklist42 - with the proportion of studies meeting each individual criterion presented in Figures 5 and 6, respectively. Most met criteria related to purpose and design. However, technical details in STRESS-DES - such as input parameters, initialisation, random sampling and execution - were often incomplete or unclear. While all studies mentioned the software or programming language used, these descriptions were minimal - for example, often mentioning the programming language and main simulation package, but not the operating system or versions used. System specifications were rarely fully described, and the requirement for pre-processing was unclear, and often marked as non-applicable. For the generic checklist, few studies addressed model uncertainties or performed sensitivity analyses, and validation was rarely discussed. Generalisability and stakeholder involvement were also seldom reported.
@@ 4.1.1 Set-up
Share code with an open licence. Half of the models were initially unlicensed, effectively preventing the code from being downloaded or run. Guidance on selecting an appropriate licence can be found in Morin et al. (2012)75 and on https://choosealicense.com/.
Link publication to a specific version of the code. In one study, updates to the code after publication meant it was unclear which version produced the reported results. Journal articles can cite the exact version of code used to generate results if it is deposited in a open science archive. These archives, such as Figshare76 or Zenodo,62 provide a digital object identifier (DOI) for citation.77 Modified versions of code can be deposited at key stages during the publication process (e.g., submission, after revision, and acceptance) to obtain a new DOI. Code repositories can also contain a changelog (e.g. https://keepachangelog.com) that can help track changes between versions.
List all dependencies and versions. Only two studies provided a complete list of required packages, and none gave all versions. Environment managers like Conda78 or renv79 are recommended for creating isolated environments and tracking dependencies. This includes recording the specific versions used, as code may break with package updates.
@@ 4.1.2 Running the model
Control randomness. Random seeds were not included in three studies. This made it impossible to determine whether differences between the results obtained from running the code and the results reported in the articles were due to variability in the simulation or issues in the code. Including a random seed is crucial for ensuring consistent results from stochastic models like DES. Tools like set.seed() in R50 or NumPy’s80 SeedSequence() and default_rng() are recommended.
@@ 4.2.2 Clarity
State run times and machine specifications. Clearly stating the expected run times and specifications of the machines used helps users understand the resource demands and assess if their machine can handle the workload, as they may not have the capacity for long or memory-intensive runs. Times can be easily recorded such as by using the time module from Python’s standard library,49 or the Sys.time() function from base R.50 It can also be helpful to record the memory usage when running the model.
@@ 4.3 Key recommendations
From the complete set of recommendations, five were identified as having the greatest impact on the reproductions:
@@ 5 Discussion
First, sharing models with an open licence. Half of the models in this research (50%, n=4/8) initially lacked a licence, similar to findings in broader healthcare DES reviews (65%, n=26/47)44 and other fields, such as computational physics (71%, n=39/55) and Science articles (34%, n=19/56).22 Open licensing is recommended by initiatives like the NHS “Goldacre review”,81 journal badge programmes,26, 63, 27 and the STARS framework.43 While private code can support internal reproducibility, it limits broader validation and reuse.
Second, maintaining an accurate and consistent record of the model parameters between the article and code was critical, as incorrect parameters often had a large impact on the observed outcomes. Third, sharing code for scenarios and sensitivity analyses - not just the base case - was vital, as reconstructing these without access to the original code was time-consuming and error-prone. These recommendations align with requirements from frameworks like STRESS-DES2 and the generic reporting checklist,42 as well as prior research identifying missing data, parameters, and scripts as major barriers to reproducibility across disciplines.22, 23, 3, 15, 18, 17, 21
The fourth and fifth recommendations involve sharing code to calculate all required model outputs and to then generate the results presented in the article (e.g. tables, figures). A lack of visualisation code or reliance on proprietary visualisation tools has similarly hindered reproduction in computational physics.17, 18 These recommendations, along with others, can help produce a reproducible analytical pipeline (RAP), wherein executing the code will run the model and produce all results from the paper. Tools like Python and R facilitate RAPs by generating figures and tables directly from code, avoiding the manual input required in drag-and-drop tools. RAPs have been highlighted as a priority by initiatives like the NHS RAP Community of Practice82 and the UK Government Analysis Function RAP Strategy.83 More widely, initiatives such as the European Commission’s exploration of reproducibility challenges in scientific research1 and a European Union project focused on RAPs for monitoring plastic pollution84 reflect the growing international emphasis on reproducibility and RAPs in diverse contexts.
@@
Some reproducibility issues are difficult to avoid, even with thorough preparation, as they may depend on future software updates or system differences. For example, one of the models used an older, now unsupported Python version (2.7.12), which may not work with some code editors/interpreters. Creating environments with an older language was straightforward in Python via Conda78 but was more challenging and not successful in R in this research. R packages are intended to be forward compatible, although this cannot be guaranteed, with newer packages breaking the code in some studies.85, 23 Similarly, one study shared a web application but this was no longer accessible remotely after its hosting site shut down. System-level code dependencies can also vary depending on the operating system used, and though they can be identified by testing the code on fresh builds, this may be beyond the scope of many projects.
@@ 5.1.1 For model developers
Common concerns about the skills9, 90 or time9, 91, 92, 90 required to prepare before sharing code can be mitigated by making a few incremental improvements throughout the project. Four of the key recommendations in this research - using correct parameters, running scenarios, producing outputs, and creating tables and figures - will already have been performed by researchers throughout the research process, and so the primary action is simply to share these complete artefacts. The fifth recommendation, adding an open licence, is a simple but essential step, enabling others to download, run, and use the code. Other recommendations from this article may require a greater time investment, though this can be minimised if considered from the outset of the project. The more recommendations that are implemented, the greater the anticipated improvement in the reproducibility of the research.
@@ 5.1.2 For peer reviewers
Regardless of whether journal or conference policies are in place, reviewers can promote reproducibility by incorporating a “reproducibility review” (Table 5.1.2). This proposed review incorporates three of the five main recommendations from this research: ensuring open licences, availability of scenario and sensitivity analysis code, and providing result-processing scripts. These checks are straightforward to identify from any provided code or artefacts. Alternatively, they could be posed as peer review questions that an author could self-certify. The remaining two recommendations—verifying consistent parameters and confirming that the model generates all outputs—were excluded, as they require a much greater time investment to investigate, including potentially needing to run the code. This review is consistent with the requirements for openly licensed and complete materials from ACM’s “Artefacts Evaluated” 26 and IEEE’s “Code Reviewed” 28 badges. The review complements other existing suggestions, such as “open scholarship review” and “longevity review” (Table 5.1.2) from Monks and Harper (2023). 44

### v2

@@ Abstract
Article word count: 8013
@@ 1.2 Addressing challenges in reproducibility
Cross-disciplinary journal initiatives to improve computational transparency and reproducibility. Initiatives Examples Mandatory code sharing Journals incentivise or require code to be shared publicly through: • Policies for mandatory code sharing.29, 30, 3, 33, 31, 32 • Badges awarded to articles with shared code.26, 27, 28 Review of shared artefacts Journals evaluate the shared artefacts with: • Policies requiring details on the language and packages, commented code, and test data/problems sufficient for replication.33 • Policies requiring open licenses and clear documentation.32 • Integration with sharing platforms which streamline submission and peer review of code within the publication workflow.31 • Badges awarded following evaluation of artefacts (e.g. checking they are complete, executable and documented).26, 28 Assess computational reproducibility Journals actively test whether results can be reproduced through: • Collaboration with independent institutions to reproduce and/or replicate a sample of published articles from the journal.32, 34 • Independent reproduction attempts during peer review, with badges awarded for successful reproductions.26, 28, 32 Accept reproductions as formal publications Journals recognise reproduction work as scholarly contribution by publishing: • Independent reproduction studies from independent groups reproducing, reusing, or extending published code.35 • Reports documenting reproduction attempts during peer review as supplementary material.26
@@ 1.3 Healthcare simulation models
This study focuses on healthcare simulation models, specifically DES. DES has a wide range of applications, including healthcare operations and system design, medical decision-making, and infectious disease modelling.5 It is the most commonly used simulation method in healthcare.38, 39, 5, 40
In recent years, efforts have been made to enhance the sharing and reuse of healthcare simulation models. The Strengthening The Reporting of Empirical Simulation Studies-DES (STRESS-DES) guidelines from Monks et al. (2019)2 were developed to facilitate replication by providing technical details necessary to recreate a model. Zhang et al. (2020)41 developed a “generic reporting checklist” focused on model quality, derived from the International Society for Pharmacoeconomics and Outcomes Research (ISPOR)–Society for Medical Decision Making (SMDM) Modelling Good Research Practices Task Force reports. In 2024, the Sharing Tools and Artefacts for Reusable Simulations (STARS) framework was introduced to provide guidance on which artefacts should be shared in repositories to facilitate model reuse.42
These guidelines and frameworks (fully detailed in Appendix A) were developed because many healthcare DES models are not shared in ways that facilitate understanding, use, and reproducibility. A 2023 review43 of 564 healthcare DES papers (2019-2022) revealed only 8.3% shared model code, although this was improving over time. Among shared models, only 60% included a README, with just 32% explaining how to run the model. Less than half provided an open licence, and only 10.6% archived their models. Across all papers, only 12.8% used reporting guidelines.43
@@
This study aims to assess the reproducibility of published healthcare DES models by examining a sample of eight studies. It will identify factors that facilitate or hinder each reproduction and will evaluate adherence to relevant frameworks, badges and reporting guidelines, to understand how these influence reproducibility. By understanding which elements are most important when reporting results in articles and sharing code, the study seeks to identify key factors that can improve reproducibility in healthcare DES models. Rather than providing a theoretical list of best practices, this empirical approach focuses on understanding what modellers actually do in practice and what specific barriers arise when attempting reproduction, enabling more targeted and actionable guidance than general principles alone. While not all studies may have been designed with computational reproducibility in mind, it is increasingly recognised as an important consideration for any research involving code,10 as it helps verify the integrity of findings. Adopting basic practices to improve reproducibility can enhance scientific rigour and is achievable without requiring advanced technical skills.
@@ 2 Methods
The study protocol (as summarised in Figure 2 and described below) was pre-registered on 20 June 2024.44 It was informed by prior reproduction studies,17, 13, 45, 46, 23 and refined through a pilot on a Python DES model by Allen et al. (2020).47
Abbreviations: STARS, Sharing Tools and Artefacts for Reusable Simulations; STRESS-DES, Strengthening The Reporting of Empirical Simulation Studies-Discrete-Event Simulation.
@@ 2.1 Sample
This study included eight models which were selected through purposive sampling. Of these, seven models were drawn from the sample of an existing systematic scoping review (2018-2022),43 whilst one additional model was identified through an informal exploratory search of earlier literature. Models were developed in Python 48 or R49, the most popular Free and Open Source Software (FOSS) languages for healthcare DES.43 Models were selected to ensure diversity across a range of factors including the health focus (e.g. healthcare condition, specific system), geographical context, and model complexity - as in Table 2.1.
The protocol initially planned to include six studies, but this was later expanded to eight during the research process. This allowed the inclusion of another “medium-sized” model 50 that was felt to better represent typical operational research applications of DES in health compared to some larger-scale models. An older model51 was also included, enabling consideration of reproducibility challenges from an older code base (approximately a decade ago), as opposed to more recent studies (with all others published within the last five years).
[…]
@@ 2.2 Reproduction
Scope. At least two team members reviewed the article and agreed on which experimental results (e.g. tables, figures, results described in the text) to reproduce. The repository was then archived on Zenodo.61
@@ 2.3 Evaluation
Articles were evaluated using two DES reporting guidelines: STRESS-DES from Monks et al. (2019),2 and the “generic reporting checklist” from Zhang et al. (2020).41 Repositories were evaluated against the STARS framework42 and criteria for journal badges from: ACM,26 National Information Standards Organisation (NISO),62 Centre for Open Science (COS),27 Institute of Electrical and Electronics Engineers (IEEE),28 and Psychological Science.63, 32 The evaluation was performed by one researcher, with consensus from a second for uncertain or unmet criteria. For reference, full guidelines and badge criteria are provided in Appendix A.
@@ 2.4 Synthesis and consolidation
Results from the reproduction, evaluation, and reflections on the facilitators and barriers encountered were shared via a Quarto website64 hosted on GitHub Pages.65 The repository was organised into a research compendium,66 with a README, conventional file organisation (e.g. data, methods, outputs), tests, and a Docker environment.67 These tests, specifically written for this study, run the model to confirm it produces consistent results, either reproducing full paper outputs or, for long-running models, executing smaller verification cases. A second researcher verified execution and reproducibility by running the model and/or tests (locally and via Docker). The final repository was archived on Zenodo,61 and results were shared with the original authors.
@@ 2.5 Citations and visualisations
Studies are not referenced by name in the results, to keep the emphasis on overarching trends rather than individual reproduction success. For transparency, full results are available, with links to the reproduction and original study repositories provided in Appendix C. For this article, plots were created using Python 3.10.14,48 Plotly 5.23.0,68 Matplotlib 3.9.269 and pandas 2.2.2,70, 71 with a full list of dependencies in the research repository.72 Other visualisations were created using Inkscape 1.3.2 (https://inkscape.org/) and Sketchpad v2022.2.21.0 (https://sketch.io/sketchpad/).
@@ 3.2 Evaluation of the repository
The repositories were also evaluated against journal badge criteria, with results for the ACM badges summarised here (Table 3.2) and the full evaluation in Appendix B. Only one repository met the standards for the “Artefacts Available” badge. None satisfied “Artefacts Functional” requirements, typically due to incomplete code (missing scenarios, outputs, etc.). Whilst those evaluations used the original artefacts as shared by the study authors, the “Results Reproduced” assessment was based on reproduction within a reasonable time allowing minimal troubleshooting, as per ACM criteria. This differs from the extended reproduction attempts described above, which allowed many hours of extensive troubleshooting. Under the stricter ACM criteria, only one study was eligible for “Results Reproduced”.
@@
Abbreviations: DES, discrete-event simulation; STARS, Sharing Tools and Artefacts for Reusable Simulations.
[…]
@@ 3.3 Evaluation of the article
Studies were evaluated against STRESS-DES2 and the generic reporting checklist41 - with the proportion of studies meeting each individual criterion presented in Figures 5 and 6, respectively. Most met criteria related to purpose and design. However, technical details in STRESS-DES - such as input parameters, initialisation, random sampling and execution - were often incomplete or unclear. While all studies mentioned the software or programming language used, these descriptions were minimal - for example, often mentioning the programming language and main simulation package, but not the operating system or versions used. System specifications were rarely fully described, and the requirement for pre-processing was unclear, and often marked as non-applicable. For the generic checklist, few studies addressed model uncertainties or performed sensitivity analyses, and validation was rarely discussed. Generalisability and stakeholder involvement were also seldom reported.
@@
Abbreviations: DES, discrete-event simulation; STRESS-DES, Strengthening The Reporting of Empirical Simulation Studies - Discrete-Event Simulation.
Abbreviations: DES, discrete-event simulation.
@@ 4.1.1 Set-up
Share code with an open licence. Half of the models were initially unlicensed, effectively preventing the code from being downloaded or run. Guidance on selecting an appropriate licence can be found in Morin et al. (2012)73 and on https://choosealicense.com/.
Link publication to a specific version of the code. In one study, updates to the code after publication meant it was unclear which version produced the reported results. Journal articles can cite the exact version of code used to generate results if it is deposited in a open science archive. These archives, such as Figshare74 or Zenodo,61 provide a digital object identifier (DOI) for citation.75 Modified versions of code can be deposited at key stages during the publication process (e.g., submission, after revision, and acceptance) to obtain a new DOI. Code repositories can also contain a changelog (e.g. https://keepachangelog.com) that can help track changes between versions.
List all dependencies and versions. Only two studies provided a complete list of required packages, and none gave all versions. Environment managers like Conda76 or renv77 are recommended for creating isolated environments and tracking dependencies. This includes recording the specific versions used, as code may break with package updates.
@@ 4.1.2 Running the model
Control randomness. Random seeds were not included in three studies. This made it impossible to determine whether differences between the results obtained from running the code and the results reported in the articles were due to variability in the simulation or issues in the code. Including a random seed is crucial for ensuring consistent results from stochastic models like DES. Tools like set.seed() in R49 or NumPy’s78 SeedSequence() and default_rng() are recommended.
@@ 4.2.2 Clarity
State run times and machine specifications. Clearly stating the expected run times and specifications of the machines used helps users understand the resource demands and assess if their machine can handle the workload, as they may not have the capacity for long or memory-intensive runs. Times can be easily recorded such as by using the time module from Python’s standard library,48 or the Sys.time() function from base R.49 It can also be helpful to record the memory usage when running the model.
@@ 4.3 Key recommendations
From the complete set of recommendations, five were identified as having the greatest impact on the reproductions, based on both fundamental enablers of reproduction and observed patterns of what consistently created barriers.
@@ 5 Discussion
First, sharing models with an open licence. Half of the models in this research (50%, n=4/8) initially lacked a licence, similar to findings in broader healthcare DES reviews (65%, n=26/47)43 and other fields, such as computational physics (71%, n=39/55) and Science articles (34%, n=19/56).22 Open licensing is recommended by initiatives like the NHS “Goldacre review”,79 journal badge programmes,26, 62, 27 and the STARS framework.42 While private code can support internal reproducibility, it limits broader validation and reuse.
Second, maintaining an accurate and consistent record of the model parameters between the article and code was critical, as incorrect parameters often had a large impact on the observed outcomes. Third, sharing code for scenarios and sensitivity analyses - not just the base case - was vital, as reconstructing these without access to the original code was time-consuming and error-prone. These recommendations align with requirements from frameworks like STRESS-DES2 and the generic reporting checklist,41 as well as prior research identifying missing data, parameters, and scripts as major barriers to reproducibility across disciplines.22, 23, 3, 15, 18, 17, 21
The fourth and fifth recommendations involve sharing code to calculate all required model outputs and to then generate the results presented in the article (e.g. tables, figures). A lack of visualisation code or reliance on proprietary visualisation tools has similarly hindered reproduction in computational physics.17, 18 These recommendations, along with others, can help produce a reproducible analytical pipeline (RAP), wherein executing the code will run the model and produce all results from the paper. Tools like Python and R facilitate RAPs by generating figures and tables directly from code, avoiding the manual input required in drag-and-drop tools. RAPs have been highlighted as a priority by initiatives like the NHS RAP Community of Practice80 and the UK Government Analysis Function RAP Strategy.81 More widely, initiatives such as the European Commission’s exploration of reproducibility challenges in scientific research1 and a European Union project focused on RAPs for monitoring plastic pollution82 reflect the growing international emphasis on reproducibility and RAPs in diverse contexts.
@@
Some reproducibility issues are difficult to avoid, even with thorough preparation, as they may depend on future software updates or system differences. For example, one of the models used an older, now unsupported Python version (2.7.12), which may not work with some code editors/interpreters. Creating environments with an older language was straightforward in Python via Conda76 but was more challenging and not successful in R in this research. R packages are intended to be forward compatible, although this cannot be guaranteed, with newer packages breaking the code in some studies.83, 23 Similarly, one study shared a web application but this was no longer accessible remotely after its hosting site shut down. System-level code dependencies can also vary depending on the operating system used, and though they can be identified by testing the code on fresh builds, this may be beyond the scope of many projects.
@@ 5.1.1 For model developers
Common concerns about the skills9, 88 or time9, 89, 90, 88 required to prepare before sharing code can be mitigated by making a few incremental improvements throughout the project. Four of the key recommendations in this research - using correct parameters, running scenarios, producing outputs, and creating tables and figures - will already have been performed by researchers throughout the research process, and so the primary action is simply to share these complete artefacts. The fifth recommendation, adding an open licence, is a simple but essential step, enabling others to download, run, and use the code. Other recommendations from this article may require a greater time investment, though this can be minimised if considered from the outset of the project. The more recommendations that are implemented, the greater the anticipated improvement in the reproducibility of the research.
@@ 5.1.2 For peer reviewers
Regardless of whether journal or conference policies are in place, reviewers can promote reproducibility by incorporating a “reproducibility review” (Table 5.1.2). This proposed review incorporates three of the five main recommendations from this research: ensuring open licences, availability of scenario and sensitivity analysis code, and providing result-processing scripts. These checks are straightforward to identify from any provided code or artefacts. Alternatively, they could be posed as peer review questions that an author could self-certify. The remaining two recommendations—verifying consistent parameters and confirming that the model generates all outputs—were excluded, as they require a much greater time investment to investigate, including potentially needing to run the code. This review is consistent with the requirements for openly licensed and complete materials from ACM’s “Artefacts Evaluated” 26 and IEEE’s “Code Reviewed” 28 badges. The review complements other existing suggestions, such as “open scholarship review” and “longevity review” (Table 5.1.2) from Monks and Harper (2023). 43
[…]